OpenStudy (precal):

ln(x^5)-ln(x^4) double checking my solution

OpenStudy (anonymous):

on a sunday????

OpenStudy (anonymous):

ln(x^5)-ln(x^4)=ln(x^5/x^4)=ln(x)

OpenStudy (precal):